OpenStudy (anonymous):

Slope intercept form? Y = -3x + 4 and -3x + 9y = 10

OpenStudy (anonymous):

@blackops2luvr

OpenStudy (samanthagreer):

Slope-intercept form is: y = mx + b where m is the slope and b is the y-intercept. Rearrange -11x + 9y = -12 to look like y = mx + b. That is, isolate y on the left side and take everything else to the right. Simplify. Identify m and b.

OpenStudy (anonymous):

-3x + 9y = 10 in slope intercept form is Y = -3x + 10 ?

OpenStudy (anonymous):

well the first equation:Y = -3x + 4 is in slope intercept, but -3x + 9y = 10 is not.

OpenStudy (anonymous):

Okay, to put -3x + 9y = 10 in slope intercept form would it be Y = -3x + 10 ?

OpenStudy (anonymous):

close, but you forgot that it is 9y.

OpenStudy (anonymous):

first part we need to get the y by itself.

OpenStudy (anonymous):

How would it be put into slope intercept form then?

OpenStudy (anonymous):

we need to get y alone, so we need to get the x away from it.

OpenStudy (anonymous):

since it's an equals statement we do this by adding the x to both sides like this: